OpenStudy (anonymous):

how do i solve square root of a to 12th power

OpenStudy (anonymous):

\[=(\sqrt{a})^{12}=(a^{1/2})^{12}=a^{(1/2)*12}=a^6\]

OpenStudy (anonymous):

how do you get the 1/2?

OpenStudy (anonymous):

The square root of any number is the same thing as taking that number to the power one-half

OpenStudy (anonymous):

The cube root is power of one-third, etc

OpenStudy (anonymous):

oh thank you so much this makes so much sense

OpenStudy (anonymous):

So if we square root a number, a, and then take it to the power of two (i.e., square it) we are left with a. This is the same as:\[\sqrt{a}*\sqrt{a}=a^{1/2}*a^{1/2}=a^1=a\]

OpenStudy (anonymous):

this example is great thank you